OpenStudy (anonymous):

Solve K =1/3 (b – 2a) for b.

OpenStudy (anonymous):

b= 3K + 2a

jimthompson5910 (jim_thompson5910):

k =1/3 (b – 2a) 3k =b – 2a 3k+2a =b b = 3k+2a